### v1

@@ IV Thermodynamic curvature via $P$-$V$ criticality
In this section, we focus on the relationship between Ruppeiner geometry (with normalized scalar curvature) and the $P$-$V$ phase transition of the spherically symmetric Hayward-AdS black hole. In thermodynamic information geometry, the rarity of a fluctuation between two thermodynamic states is indicative of their greater separation. Consequently, the line element Wei:2019uqg ; Wei:2019yvs
$dl^{2}$ $=$ $\dfrac{C_{V}}{T^{2}}dT^{2}-\dfrac{\left(\partial_{V}P\right)_{T}}{T}dV^{2}\,$, (30)
contains information regarding the effective interaction that exists between two microscopic fluctuation states. For a specific fluid system, the curvature scalar derived from Eq.(30) serves its microstructure interactions. Notably, the entropy within this information geometry framework, known as Ruppeiner geometry, constitutes the thermodynamic potential.
For the Hayward-AdS black hole, the constant-volume heat capacity $C_{V}=T(\partial_{T}S)_{V}$ equals zero. This is because the entropy $S$ of a black hole is a function of the horizon radius $r_{+}$ $i.e.$, $S(r_{+})$. Consequently, both the line element (30) and the curvature scalar $R$ exhibit divergence. To avoid this divergence, a new normalized curvature scalar $R_{N}$ is defined as Wei:2019uqg ; Wei:2019yvs
$R_{\rm N}=RC_{V}\,.$ (31)
This normalized curvature scalar is more penetrating to investigate the microscopic properties of black holes. After straightforward calculations, the normalized scalar curvature of the Hayward-AdS black hole expressed in terms of reduced parameters as follows
@@ V Conclusions and Discussion
Furthermore, we use Ruppeiner geometry to investigate the microstructures of Hayward-AdS black holes—an approach that also enables probing of their thermodynamic fluctuations. We first calculated the normalized scalar curvature ($R_{\rm N}$), a quantity whose sign encodes information about the dominant microscopic interactions and which is linked to the correlation length near the critical point. Solving for $R_{\rm N}$ allowed us to identify two key temperatures: $\widetilde{T}_{\rm div}$ (the temperature at which $R_{\rm N}$ diverges) and $\widetilde{T}_{0}$ (the temperature at which $R_{\rm N}=0$). Notably, within the coexisting phase volume of small and large black holes, two such divergence temperatures exist, and these temperatures coincide exactly at the critical point. Our analysis confirms the existence of $\widetilde{T}_{0}$ and reveals a clear transition in dominant microscopic interactions: near the critical point, attractive interactions prevail, while repulsive interactions become dominant at lower temperatures. This interaction structure is universal across most AdS black holes, and our results thus provide preliminary but valuable insights into the microstructural properties of higher-dimensional Hayward-AdS black holes—along with direct understanding of the interaction nature of their microscopic constituents.
@@ Acknowledgements
This work is supported by National Natural Science Foundation of China (NSFC) under grant No. 12465012, Kashi University high-level talent research start-up fund project under grant No. 022024002, and Tianchi Talented Young Doctors Program of Xinjiang Uyghur Autonomous Region.

### v2

@@ IV Thermodynamic curvature via $P$-$V$ criticality
In this section, we focus on the relationship between Ruppeiner geometry (with normalized scalar curvature) and the $P$-$V$ phase transition of the spherically symmetric Hayward-AdS black hole. Ruppeiner geometry is a conceptual framework rooted in Riemannian geometry and the fluctuation theory of equilibrium thermodynamics, where entropy plays a central role. Its line element, expressed in terms of entropy, takes the form Ruppeinerb2008 ; Ruppeiner95
${\rm d}{l^{2}}=g_{\mu\nu}{\rm d}{x^{\mu}}{\rm d}{x^{\nu}}\,$, (30)
where ${\rm d}{l^{2}}$ denotes the distance between two adjacent fluctuation states. The fluctuation coordinates $x=\left({U,V}\right)$ correspond to internal energy $U$ and volume $V$, and the metric tensor is given by ${g_{\mu\nu}}=-{\partial_{\mu,\nu}}S$. For black hole systems, the first law of thermodynamics reads Ruppeinerb2008 ; Ruppeiner95
${\rm d}S=\frac{1}{T}{\rm d}U+\frac{P}{T}{\rm d}V\,.$ (31)
Comparing Eqs. (30) and (31) and identifying the conjugate quantities of $x$ as ${y_{\mu}}=\partial S/\partial{x^{\mu}}$, leads to the following relations
${{\rm d}}\left({\frac{1}{T}}\right)={\left({\frac{{{\partial^{2}}S}}{{\partial{U^{2}}}}}\right)_{V}}{{\rm d}}U+\left({\frac{{{\partial^{2}}S}}{{\partial U\partial V}}}\right){{\rm d}}V\,$, (32a)
${{\rm d}}\left({\frac{P}{T}}\right)={\left({\frac{{{\partial^{2}}S}}{{\partial{V^{2}}}}}\right)_{U}}{{\rm d}}V+\left({\frac{{{\partial^{2}}S}}{{\partial U\partial V}}}\right){{\rm d}}U\,.$ (32b)
The line element in Eq. (30) can thus be rewritten as
${{\rm d}}{l^{2}}$ $=-{{\rm d}}\left({\frac{1}{T}}\right){{\rm d}}U-{{\rm d}}\left({\frac{P}{T}}\right){{\rm d}}V$
$=\frac{1}{{{T^{2}}}}{{\rm d}}T{{\rm d}}U-\frac{1}{T}{{\rm d}}P{{\rm d}}V+\frac{P}{{{T^{2}}}}{{\rm d}}T{{\rm d}}V\,.$ (33)
[…]
${\rm d}l^{2}$ $=$ $\dfrac{C_{V}}{T^{2}}{\rm d}T^{2}-\dfrac{\left(\partial_{V}P\right)_{T}}{T}{\rm d}V^{2}\,$, (34)
which encodes information about the effective interactions between microscopic fluctuation states. For a given fluid system, the curvature scalar derived from Eq. (34) characterizes its microstructural interactions. Notably, in Ruppeiner geometry–an information geometry framework where entropy serves as the thermodynamic potential–the black hole entropy $S$ is a function of the horizon radius $r_{+}$ $i.e.$, $S(r_{+})$. This implies that the constant-volume heat capacity $C_{V}=T(\partial_{T}S)_{V}$ vanishes, leading to divergences in both the line element (34) and the curvature scalar $R$.
To resolve this divergence, a normalized curvature scalar $R_{N}$ is defined as Wei:2019uqg ; Wei:2019yvs
$R_{\rm N}$ $=$ $RC_{V}$ (35)
$=$ $\frac{{\left({{\partial_{V}}P}\right)_{T}^{2}-{T^{2}}{{\left({{\partial_{T,V}}P}\right)}^{2}}+2{T^{2}}{{\left({{\partial_{V}}P}\right)}_{T}}\left({{\partial_{T,T,V}}P}\right)}}{{2\left({{\partial_{V}}P}\right)_{T}^{2}}}\,.$
This normalized curvature scalar provides a more robust tool for investigating the microscopic properties of black holes.
After straightforward calculations, the normalized scalar curvature of the Hayward-AdS black hole expressed in terms of reduced parameters as follows
@@ V Conclusions and Discussion
Furthermore, we use Ruppeiner geometry to investigate the microstructures of Hayward-AdS black holes–an approach that also enables probing of their thermodynamic fluctuations. We first calculated the normalized scalar curvature ($R_{\rm N}$), a quantity whose sign encodes information about the dominant microscopic interactions and which is linked to the correlation length near the critical point. Solving for $R_{\rm N}$ allowed us to identify two key temperatures: $\widetilde{T}_{\rm div}$ (the temperature at which $R_{\rm N}$ diverges) and $\widetilde{T}_{0}$ (the temperature at which $R_{\rm N}=0$). Notably, within the coexisting phase volume of small and large black holes, two such divergence temperatures exist, and these temperatures coincide exactly at the critical point. Our analysis confirms the existence of $\widetilde{T}_{0}$ and reveals a clear transition in dominant microscopic interactions: near the critical point, attractive interactions prevail, while repulsive interactions become dominant at lower temperatures. This interaction structure is universal across most AdS black holes, and our results thus provide preliminary but valuable insights into the microstructural properties of higher-dimensional Hayward-AdS black holes along with direct understanding of the interaction nature of their microscopic constituents.
@@ Acknowledgements
This work is supported by National Natural Science Foundation of China (Grant No. 12465012), Kashi University high-level talent research start-up fund project (Grant No. 022024002).